(2y+3)^2-7=0 using the square root property.
Sqrt (3-x) ) -7=0
Sqrt (x+5) )= 7-x
Simplify 2 sqrt (75))- 5 sqrt (27))- sqrt (48))

Answer by edjones(8007)

(2y+3)^2-7=0 using the square root property.
(2y+3)^2=7
2y+3=+-sqrt(7)
2y=-3+-sqrt(7)
y=(-3+-sqrt(7))/2
.
Ed